# Quantum phase transitions in the alternating XY chain with three-site interaction

### 02 Jan 2024 v1

@@ Abstract
We investigate the quantum phase transition in the alternating XY chain with the XZY+YZX type of three-spin interactions. We present the exact solution derived by means of the Jordan-Wigner transformation and study the average magnetization, spin correlations, and von Neumann entropy to establish the phase diagram. By examining the nearest-neighbor transverse spin correlation, we probe that in the phase with zero magnetization (ZM), the spins within a supercell tend to point to opposite directions of the external field, but between the nearest-neighbor supercells are distributed randomly. This is the reason for the magnetization vanishing in the ZM phase under the alternating hopping interaction. In addition, we also investigate the influence of the three-site interaction, and find that the ZM phase is absent as the strength of the three-site interaction increases. Our findings shed light on the complex behavior of the alternating XY chain and provide valuable insights for future studies.
@@ I Introduction
In this paper, we consider the alternating XY chain Perk et al. (1975); Siskens et al. (1975b) with XZX+YZY type of three-site interaction in the transverse field [see Fig. 1]. We show that this model is exactly solvable and obtain the ground state phase diagram by calculating the average magnetization and spin correlations. By examining the nearest-neighbor transverse spin correlation, we probe that in the phase with zero magnetization (ZM), spins within a supercell point to the opposite direction along the external field, and spins between supercells are not coupled. Additionally, by investigating the von Neumann entropy, we notice that the von Neumann entropy in the phase with zero magnetization has certainly large value while is different from the cases in the paramagnetic phase. This exactly explains the reason for the magnetization vanishing in the so-called ZM phase. The ZM phase induced by the alternating interaction thus should not be classified as the PM phase. In addition, we also discuss the influence of the three-site interaction, and find that the zero magnetization phase is absent as the strength of the three-site interaction increases, instead of the appearance of two ferromagnetic phases.
@@ III Order parameters
as a function of the external field $h$ for the three typical cases as shown in Fig. 3 and 4. From Figs. (a) and (b), it can be observed that the average magnetization $M_{z}$ exhibits plateaus (remain constant with the increase of $h$) at phases I, V, and IV, corresponding to $M_{z}=-1$ in phase I ($h<h_{c2}^{-}$), $M_{z}\approx 0$ in phase V, and $M_{z}=1$ in phase IV. It is evident that phases I and IV exhibit the paramagnetic behavior here. Additionally, the first-order derivatives $\frac{dM_{z}}{dh}$ of the average magnetization show the nonanalytic singularities at the critical points between each phase.
@@ IV Quantum entanglement
In Fig. 9, we display the von Neumann entropy as a function of the size $L$ for the system within each phase. For simplicity, we only show the results of the system with $\Omega=0.4$, for which phases V, VI, and VII appear simultaneously. Since the definition of von Neumann entropy is based on the local block in the finite chain, the values of von Neumann entropy are observed to be different for the odd and even $L$. For each case shown in Fig. 9, we can see that the von Neumann entropy has a small value approaching zero in phases I and IV. This indicates that the states in phases I and IV are deterministic, consistent with the behavior of the PM phase where all spins point to the direction of the external field. However, the von Neumann entropy in phase V has certainly large value, suggesting its difference from that in the PM phase. The result is easy to explain that the spins between the supercells are not strongly coupled, so the spins between the supercells are distributed randomly.
@@ Appendix D Case of the dimerized lattice
Now, we show more evidence that the spins within a supercell point to the opposite direction along the external field in the dimerized limits. In Fig. 12, we show the average transverse spin-spin correlation functions $TC_{\mathrm{odd}}$ and $TC_{\mathrm{even}}$ as functions of the external field $h$ for $\alpha=\frac{J_{0}-J_{1}}{J_{0}+J_{1}}=0.2$. It can be seen that, as compared to the results of $\alpha=0.5$ (see Fig. 8), the correlation functions $TC_{\mathrm{odd}}$ and $TC_{\mathrm{even}}$ show the clearer plateaus in the ZM phase, corresponding to $TC_{\mathrm{odd}}\approx-0.998\approx-1$ and $TC_{\mathrm{even}}\approx 0$. The results indicate that in the dimerized limit, the spins within a supercell construct an anti-paramagentic cluster in the ZM phase, and the spins between the nearest-neighbor supercells have no correlations, corresponding to the random distribution of the supercells.

### 04 Jan 2024 v2

@@ Abstract
We investigate the quantum phase transition in the alternating XY chain with the XZX+YZY type of three-spin interactions. We present the exact solution derived by means of the Jordan-Wigner transformation and study the average magnetization, spin correlations, and von Neumann entropy to establish the phase diagram. The phase diagram consists of the ferromagnetic phases, the paramagnetic phases, and the phase with weak magnetization (WM). By examining the nearest-neighbor transverse spin correlation, we probe that in the WM phase, the spins within a supercell generate a cluster with a small total spin, but between the nearest-neighbor supercells are distributed randomly. Especially for the dimerized limit case, the spins within a supercell tend to point to opposite directions of the transverse field. In addition, we also investigate the influence of the three-site interaction, and find that the WM phase is absent as the strength of the three-site interaction increases. Our findings shed light on the complex behavior of the alternating XY chain and provide valuable insights for future studies.
@@ I Introduction
In this paper, we consider the alternating XY chain Perk et al. (1975); Siskens et al. (1975b) with XZX+YZY type of three-site interaction in the transverse field [see Fig. 1], which is exactly solvable. A similar model was studied by Zvyagin in Ref. Zvyagin, 2009 before, but the phase diagram of the current model is still unclear and requires further investigation to understand the properties of each phase. By calculating the average magnetization, spin correlations, and von Neumann entropy, we obtain the ground state phase diagram, consisting of the ferromagnetic phases, the paramagnetic phases, and the phase with weak magnetization (WM). By examining the nearest-neighbor transverse spin correlation, we find that in the WM phase, spins within a supercell generate a cluster with a small total spin, but between the supercells are distributed randomly. The total spin of a cluster in a supercell is dependent on the dimerized structure, denoted by the ratio $\frac{J_{0}-J_{1}}{J_{0}+J_{1}}$. In particular, we observe that for the dimerized limit case, the spins within a supercell tend to point to opposite directions of the external field. This is the reason for the weak magnetization (or even zero magnetization in the fully dimerized case) in the WM phase. Moreover, by investigating the von Neumann entropy, we notice that the von Neumann entropy in the WM phase has certainly large value which is different from the cases in the paramagnetic phase. The WM phase induced by the alternating interaction thus should not be classified as the PM phase. In addition, we also discuss the influence of the three-site interaction, and find that the WM phase is absent as the strength of the three-site interaction increases, instead of the appearance of two ferromagnetic phases.
@@ III Order parameters
as a function of the external field $h$ for the three typical cases as shown in Fig. 3 and 4. From Figs. (a) and (b), it can be observed that the average magnetization $M_{z}$ exhibits plateaus (remain constant with the increase of $h$) at phases I, V, and IV, corresponding to $M_{z}=-1$ in phase I ($h<h_{c2}^{-}$), $M_{z}=1$ in phase IV, and a finite value approaching zero ($M_{z}\approx 0$) in phase V. It is evident that phases I and IV exhibit the paramagnetic behavior here. Additionally, the first-order derivatives $\frac{dM_{z}}{dh}$ of the average magnetization show the nonanalytic singularities at the critical points between each phase.
$\rho^{x}$ $=$ $\lim_{r\rightarrow\infty}C_{r}^{x},$ (19)
$\rho^{y}$ $=$ $\lim_{r\rightarrow\infty}C_{r}^{y}.$ (20)
@@ IV Quantum entanglement
In Fig. 9, we display the von Neumann entropy as a function of the size $L$ for the system within each phase. For simplicity, we only show the results of the system with $\Omega=0.4$, for which phases V, VI, and VII appear simultaneously. Since the definition of von Neumann entropy is based on the local block in the finite chain, the values of von Neumann entropy are observed to be different for the odd and even $L$. For each case shown in Fig. 9, we can see that the von Neumann entropy has a small value approaching zero in phases I and IV. This indicates that the states in phases I and IV are deterministic, consistent with the behavior of the PM phase where all spins point to the direction of the external field. However, the von Neumann entropy in phase V has certainly large value, suggesting its difference from that in the PM phase. The result is easy to explain that the spins between the supercells are distributed randomly in phase V so that the system possesses large von Neumann entropy.
@@ Appendix D Case of the dimerized lattice
Now, we show more evidence that the spins within a supercell point to the opposite direction along the external field in the dimerized limits. In Fig. 12, we show the average transverse spin-spin correlation functions $TC_{\mathrm{odd}}$ and $TC_{\mathrm{even}}$ as functions of the external field $h$ for $\alpha=\frac{J_{0}-J_{1}}{J_{0}+J_{1}}=0.1$. It can be seen that, as compared to the results of $\alpha=0.5$ (see Fig. 8), the correlation functions $TC_{\mathrm{odd}}$ and $TC_{\mathrm{even}}$ show the clearer plateaus in phase V, corresponding to $TC_{\mathrm{odd}}\approx-0.998\approx-1$ and $TC_{\mathrm{even}}\approx 0$. The results indicate that in the dimerized limit, the spins within a supercell construct an anti-paramagentic cluster in phase V, and the spins between the nearest-neighbor supercells are not coupled, corresponding to the random distribution of the supercells.